OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty dollars. This year, he paid one hundred sixty-nine dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

166 - 130 = 36 36/130 = .2769 * 100 27.69%

OpenStudy (anonymous):

(39/130) * 100 = 30 %

OpenStudy (anonymous):

its 169-130 = 39 dollars more